# DIsoN: Decentralized Isolation Networks for Out-of-Distribution Detection in Medical Imaging

### v1

@@ Abstract
Safe deployment of machine learning (ML) models in safety-critical domains such as medical imaging requires detecting inputs with characteristics not seen during training, known as out-of-distribution (OOD) detection, to prevent unreliable predictions. Effective OOD detection after deployment could benefit from access to the training data, enabling direct comparison between test samples and the training data distribution to identify differences. State-of-the-art OOD detection methods, however, either discard the training data after deployment or assume that test samples and training data are centrally stored together, an assumption that rarely holds in real-world settings. This is because shipping the training data with the deployed model is usually impossible due to the size of training databases, as well as proprietary or privacy constraints. We introduce the Isolation Network, an OOD detection framework that quantifies the difficulty of separating a target test sample from the training data by solving a binary classification task. We then propose Decentralized Isolation Networks (DIsoN), which enables the comparison of training and test data when data-sharing is impossible, by exchanging only model parameters between the remote computational nodes of training and deployment. We further extend DIsoN with class-conditioning, comparing a target sample solely with training data of its predicted class. We evaluate DIsoN on four medical imaging datasets (dermatology, chest X-ray, breast ultrasound, histopathology) across 12 OOD detection tasks. DIsoN performs favorably against existing methods while respecting data-privacy. This decentralized OOD detection framework opens the way for a new type of service that ML developers could provide along with their models: providing remote, secure utilization of their training data for OOD detection services. Code will be made available upon acceptance at: ************
11footnotetext: Equal contribution.
@@ 1 Introduction
Consider the standard setting where an organization, such as a software company, develops and trains a Machine Learning (ML) model to perform a task of interest, for example disease classification in medical images, using a training database. The model is then provided to a user (e.g. client) and deployed, for example in a hospital, to make predictions on new test data samples. Because of heterogeneity in real-world data or user-error, a deployed model may receive inputs unlike anything seen in the training data. Such inputs are called out-of-distribution (OOD) samples, in contrast to the in-distribution (ID) samples that follow the distribution of the training data. For example, OOD patterns in medical imaging can be artifacts from suboptimal acquisition or unknown diseases. Performance of ML models often degrades unexpectedly on OOD data [38]. Thus to ensure safe deployment in safety-critical applications, such as healthcare, deployment frameworks should include mechanisms for detecting OOD inputs, so that they can be flagged to human users and avoid adverse effects of using potentially wrong model predictions in the downstream workflow.
Multiple OOD detection methods have been previously developed, which can be broadly categorized into post-hoc and training-time regularization methods [37]. Post-hoc methods assume that a "primary" model has been trained and deployed to perform a task of interest, such as disease classification, and they aim to perform OOD detection without alterations to this primary model [12, 20, 21, 25, 14]. On the other hand, training-time regularization methods can be applied when training the primary model, altering its embedding space such that it will allow improved OOD detection during inference [28, 27, 5, 13, 32, 26].
Few ODD detection methods leverage direct comparison of test-samples against the training data to identify differences between them for OOD detection, such as KNN-based [30] or density-based methods [6]. These require, however, the training data to be available at the site of deployment, for direct comparison with test-samples. Therefore they are impractical in many applications where sharing and storing the training samples or their embeddings at each site of deployment is impossible due to privacy, legal, or proprietary constraints. Because of this, most other methods do not compare the test-samples against the training data [12, 19] or make indirect comparisons using auxiliary representations of training data, such as via summary statistics [20], prototypes [27, 28] or synthetic samples [7]. Such derived representations, however, do not faithfully capture all intricacies of the original training data.  Hypothesizing that direct comparison with the original training data would be useful to infer whether a test sample is OOD, this paper addresses the following question:
@@
This paper describes a novel OOD detection framework (Fig. 1) with the following key aspects: (1) We introduce Isolation Networks for OOD detection. To infer whether a new test sample is OOD, a neural network is trained to learn a binary classification boundary that separates (isolates) a test sample from the original ID training samples. The network’s convergence rate is then used as a measure of whether the target sample is OOD or similar to the training data. The intuition is that test samples with OOD patterns will be easier to separate from the training samples than ID test samples without OOD patterns. Isolation Networks draw inspiration from Isolation Forests [23], which train decision trees to isolate each training sample. To infer whether a test-sample is OOD, the number of split nodes needed to isolate it is used (c.f. Sec. 2). Isolation Networks instead train a different network for each test sample, to separate it from the training data, and measure convergence as OOD score. (2) We then introduce Decentralized Isolation Networks (DIsoN), a decentralized training framework that enables training Isolation Networks in the practical setting when the ID training data are held on a different computational node than the node of deployment, where the test samples are processed. (3) To perform class-conditional training of a DIsoN OOD detector, where the class of a test sample is predicted, which is then used to compare it only against samples in the ID training database of the same class. This reduces variability within the distribution of training samples that the model uses for comparison, making it harder to isolate ID samples that closely resemble their class than OOD samples, improving OOD detection.
@@ 2 Related Work
Federated Learning with OOD Detection. Training of DIsoN uses decentralised optimization similar to Federated Learning (FL). Using FL for OOD detection is a recent area. The few related works, such as  [39, 22], studied settings significantly different from ours: they assume the training data is decentralized across multiple computational nodes, where each node’s data follow a different distribution. Using FL, they train a model that measures the distribution-shifts between nodes. After training, the model is applied on test-samples for OOD detection. Our algorithm does not require training data from multiple distributions or multiple nodes. It assumes one node holds all ID training data and a second node holds a single test-sample, which we infer whether it is OOD via decentralised training of a binary classifier. Thus the motivation, use-case and technical challenges are distinct.
@@ 3.1 Isolation Networks
The core idea of our approach is that the difficulty of training a binary classifier to separate a single target sample $\mathbf{x}_{t}$ from the source data $\mathcal{D}_{s}$ gives an indication whether $\mathbf{x}_{t}$ is ID or OOD. Intuitively, a target sample that is ID will share characteristics with samples in $\mathcal{D}_{s}$. As a result, it is harder for a classifier to isolate (separate) the target sample and will require more update steps. On the other hand, an OOD target sample has patterns that differ from $\mathcal{D}_{s}$, making it easier to isolate and require fewer update steps during training. Fig. 2 gives a visual intuition of the idea.
Formally, we consider a neural network consisting of a feature extractor $f:\mathcal{X}\to\mathcal{Z}$, parameterized by $\theta^{f}$ and a binary classification head $h:\mathcal{Z}\to[0,1]$, with parameters $\theta^{h}$. The full network is parameterized by $\theta=(\theta^{f},\theta^{h})$. In an idealized centralized setting with full access to $\mathcal{D}_{s}$ and the target sample $\mathbf{x}_{t}$, we can train the network $h\circ f$ for the following binary classification problem: Assign label $y=0$ to all source samples $\mathbf{x}_{s}\in\mathcal{D}_{s}$and label $y=1$ to the target sample $\mathbf{x}_{t}$. Typically, $\mathcal{D}_{s}$ has a high number of samples compared to the single target sample $\mathbf{x}_{t}$, $|\mathcal{D}_{s}|\gg 1$. Therefore, directly training on this highly imbalanced setup is suboptimal, as it can lead to a trivial classifier that always predicts the majority class ($y=0$), ignoring the minority class ($y=1$). To ensure that the target sample has an impact during training, we over-sample $\mathbf{x}_{t}$ and construct mini-batches $B$ by sampling a set of source samples $B_{s}$ from $\mathcal{D}_{s}$ and we include the target sample $\mathbf{x}_{t}$ $N$-times in the mini-batch. The empirical loss for this optimization objective is:
$\mathcal{L}_{c}(\theta;B_{s},\mathbf{x}_{t},N)=\frac{1}{|B_{s}|+N}\left(\sum_{\mathbf{x}_{s}\in B_{s}}L(\theta;\mathbf{x}_{s},0)+N\cdot L(\theta;\mathbf{x}_{t},1)\right),$ (1)
where $L(\theta;x,y)$ denotes the binary cross-entropy loss. In Sec. 3.2 we compare the gradient of the centralized version $g_{c}(\theta;B_{s},\mathbf{x}_{t},N)=\nabla_{\theta}\mathcal{L}_{c}(\theta;B_{s},\mathbf{x}_{t},N)$ to that of our proposed decentralized version.
@@
Let $p^{(k)}(x)=h(f(x;\theta^{(k)}))$ be the sigmoid-score after $k$ updates. Fix a window $E_{\mathrm{stab}}\in\mathbb{N}$ and accuracy threshold $\tau\in(0,1]$. The convergence time $K$ is the smallest $k\geq E_{\mathrm{stab}}$ such that
@@
Intuitively, the conditions ensure that the classifier correctly classifies $\mathbf{x}_{t}$ for the last $E_{\mathrm{stab}}$ consecutive updates (first condition), while achieving an accuracy of at least $\tau$ on the source data (second condition). We set $E_{\mathrm{stab}}=5$ and $\tau=0.85$. Ergo, we define our OOD scoring function as $S(\mathbf{x}_{t})=K$, aligning with the convention where ID samples have higher scores.
@@ 3.2 Decentralized Isolation Networks (DIsoN)
where $g_{S}(\theta)=\frac{1}{|B_{s}|}\sum_{\mathbf{x}_{s}\in B_{s}}\nabla_{\theta}L(\theta;\mathbf{x}_{s},0)$, $g_{T}(\theta)=\nabla_{\theta}L(\theta;\mathbf{x}_{t},1)$ and $N$ the number of times $\mathbf{x}_{t}$ is oversampled in the centralized version.
@@
Prop. 3.1 shows that our decentralized training exactly replicates the centralized version when $E=1$ and aggregation weights are chosen accordingly. This theoretical equivalence motivated our design of DIsoN: it preserves the core idea of the Isolation Network, while meeting our decentralized data sharing constraints. In practice, we allow $E>1$ to reduce communication overhead, which causes the decentralized updates to deviate from the centralized version. However, as we show in Sec. 4, DIsoN achieves promising results even with the approximation. We found that $\alpha=0.8$ consistently performs well across all our experiments. In Sec. 4.2, we conduct an ablation study on this hyperparameter.
Practical Techniques: Augmentation and Normalization. One challenge in DIsoN is to avoid overfitting the target sample $\mathbf{x}_{t}$. The model may rapidly memorize superficial features of $\mathbf{x}_{t}$, making the isolation task trivial regardless whether the sample is OOD. To prevent this, we apply stochastic data augmentations (e.g. random crops, horizontal flips), which act as a regularizer and encourage the model to learn invariant features, so that the separation is based on semantic characteristics. Furthermore, in DIsoN, we use Instance Normalization (IN) [31] instead of the widely used Batch Normalization (BN) [15] layers for feature normalization. BN relies on batch-level statistics, which is not suitable for our single-sample TN. IN instead normalizes each feature map per sample and therefore fits better for our decentralized, single target sample scenario.
@@ 4 Experiments & Results
Datasets. We evaluate DIsoN on four publicly available medical imaging benchmark datasets covering dermatology, breast ultrasound, chest X-ray, and histopathology. The first three benchmarks use images with non-diagnostic artifacts as OOD samples (e.g., rulers, pacemakers, annotations), while histopathology focuses on semantic and covariate shifts across domains. Example images are shown in Fig.3.
@@
Table 1: OOD detection performance evaluated on three OOD datasets: Chest X-Ray, Dermatology, and Breast Ultrasound. $\downarrow$ means smaller is better and $\uparrow$ means larger is better. Bold numbers highlight the best results, second best results are underlined.
Method Chest X-Ray Skin Lesion Breast Ultrasound Average
@@
MSP 100.00 63.39 100.00 61.52 100.00 54.30 100.00 59.74
MDS 90.00 55.53 60.56 72.68 73.21 60.82 74.59 63.01
fDBD 83.10 67.87 96.48 61.90 91.07 59.12 90.22 62.96
ViM 78.60 67.71 71.83 70.56 69.64 56.50 73.36 64.92
iForest 92.30 52.50 89.40 49.30 94.60 49.50 92.10 50.43
CIDER 68.50 77.68 60.56 83.43 85.70 61.00 71.59 74.04
PALM 79.20 69.89 61.97 76.21 76.79 59.59 72.65 68.56
CC-DIsoN 60.52 83.94 40.14 87.96 67.86 66.58 56.17 79.49
Figure 3: Examples of data. X-Ray: ID X-ray vs. OOD scan with pacemaker. Dermatology: ID lesion vs. OOD image with ruler. Ultrasound: ID artifact-free ultrasound vs. OOD scan containing annotations. Histopathology: ID mitotic-cell patch vs. near-OOD patch with different cancer type and far-OOD patch with different staining.
Training Details. We use a ResNet18 [11] with Instance Normalization (as per Sec.3.2) pre-trained on the dataset-specific task as initialization for DIsoN. DIsoN is trained with Adam (lr=0.001 for dermatology and ultrasound; 0.003 for X-ray). For histopathology SGD with momentum (lr=0.01, momentum=0.9) is used (since [10] suggests pretraining with SGD). Local iterations per communication round are chosen to approximately match one epoch on the training data. We use standard augmentations (e.g. random cropping, rotation, color-jitter) and the aggregation weight is fixed to $\alpha=0.8$ (see Sec.  4.2 for effect of $\alpha$). Experiments were run on an NVIDIA RTX A5000. More training details and hyperparameters are provided in the Appendix. OOD detection performance is evaluated with two metrics: (i) area under the receiver operating characteristics curve (AUROC) (ii) the false positive rate at 95% true positive rate (FPR95).
Baselines. We compare our method against state-of-the-art OOD detection methods from the two main categories: post-hoc and training-time regularization methods. The post-hoc methods include: MSP [12], MDS [20], fDBD [24], ViM [33], Deep iForest [35]. For training-time regularization methods, we evaluate the recent methods CIDER [28] and PALM [27] (both using contrastive learning). We use MDS for OOD scoring on their learned feature representations.
4.1 Evaluation on Medical OOD Benchmarks
Dermatology, Chest X-Ray and Breast Ultrasound.
@@ 4.1 Evaluation on Medical OOD Benchmarks
MSP 91.6/58.2 86.1/57.5 94.8/54.2 94.0/59.0 93.2/54.1 91.2/51.5 96.4/52.1 92.5/55.2 62.5/79.1 64.9/83.1 63.8/81.1
MDS 76.5/70.2 78.5/67.2 78.5/65.6 83.7/64.6 84.5/62.4 86.8/60.4 89.2/51.9 82.5/63.2 35.9/93.3 41.8/87.5 38.9/90.4
fDBD 82.5/62.4 82.5/57.2 86.1/57.7 86.8/62.2 89.6/58.8 92.4/53.6 91.2/54.5 87.3/58.1 61.8/82.7 64.5/82.7 63.1/82.7
ViM 71.7/71.4 72.9/69.7 75.7/64.8 83.3/65.0 86.5/62.8 87.2/61.1 95.6/53.6 81.8/64.1 26.7/94.4 32.3/91.3 29.5/92.9
iForest 99.2/33.6 99.2/40.8 98.8/38.6 98.8/36.1 98.8/37.3 99.2/40.9 96.0/45.0 98.6/38.9 99.6/20.1 99.2/27.3 99.4/23.7
CIDER 90.0/74.5 89.6/74.7 93.6/61.9 89.6/61.8 95.6/55.2 96.8/52.1 94.0/62.6 92.8/63.2 80.1/81.0 27.9/92.0 54.0/86.5
PALM 76.9/75.9 88.5/70.3 83.3/68.6 93.2/63.0 89.6/64.0 97.2/46.9 95.6/63.1 89.2/64.5 1.6/98.9 0.0/99.9 0.8/99.4
CC-DIsoN 78.8/75.5 64.0/79.3 78.8/72.0 87.2/61.2 86.8/64.5 79.2/70.8 91.3/59.6 80.9/69.0 5.2/98.2 4.8/98.3 5.0/98.3

### v2

@@ Abstract
Safe deployment of machine learning (ML) models in safety-critical domains such as medical imaging requires detecting inputs with characteristics not seen during training, known as out-of-distribution (OOD) detection, to prevent unreliable predictions. Effective OOD detection after deployment could benefit from access to the training data, enabling direct comparison between test samples and the training data distribution to identify differences. State-of-the-art OOD detection methods, however, either discard the training data after deployment or assume that test samples and training data are centrally stored together, an assumption that rarely holds in real-world settings. This is because shipping the training data with the deployed model is usually impossible due to the size of training databases, as well as proprietary or privacy constraints. We introduce the Isolation Network, an OOD detection framework that quantifies the difficulty of separating a target test sample from the training data by solving a binary classification task. We then propose Decentralized Isolation Networks (DIsoN), which enables the comparison of training and test data when data-sharing is impossible, by exchanging only model parameters between the remote computational nodes of training and deployment. We further extend DIsoN with class-conditioning, comparing a target sample solely with training data of its predicted class. We evaluate DIsoN on four medical imaging datasets (dermatology, chest X-ray, breast ultrasound, histopathology) across 12 OOD detection tasks. DIsoN performs favorably against existing methods while respecting data-privacy. This decentralized OOD detection framework opens the way for a new type of service that ML developers could provide along with their models: providing remote, secure utilization of their training data for OOD detection services. Code available at: https://github.com/FelixWag/DIsoN
11footnotetext: Equal second-author contribution.
@@ 1 Introduction
Consider the standard setting where an organization, such as a software company, develops and trains a Machine Learning (ML) model to perform a task of interest, for example disease classification in medical images, using a training database. The organization then provides the model to a user (e.g., a client) and deploys it, for example in a hospital, to make predictions on new test samples. Because of heterogeneity in real-world data or user-error, a deployed model may receive inputs unlike anything seen in the training data. Such inputs are called out-of-distribution (OOD) samples, in contrast to the in-distribution (ID) samples that follow the distribution of the training data. For example, OOD patterns in medical imaging can be artifacts from suboptimal acquisition or unknown diseases. Performance of ML models often degrades unexpectedly on OOD data [43]. Thus, to ensure safe deployment in safety-critical applications such as healthcare, deployment frameworks should include mechanisms to detect OOD inputs, so that they can be flagged to human users and avoid adverse effects of using potentially wrong model predictions in the downstream workflow.
Multiple OOD detection methods have been previously developed, which can be broadly categorized into post-hoc and training-time regularization methods [42]. Post-hoc methods assume that a “primary” model has been trained and deployed to perform a task of interest, such as disease classification, and they aim to perform OOD detection without alterations to this primary model [13, 22, 23, 27, 15]. On the other hand, training-time regularization methods alter the embedding space of the primary model during its training, to allow improved OOD detection during inference [30, 29, 5, 14, 36, 28].
Few OOD detection methods leverage direct comparison of test-samples against the training data to identify differences between them for OOD detection, such as KNN-based [34] or density-based methods [6]. These require, however, the training data to be available at the site of deployment, for direct comparison with test-samples. Therefore they are impractical in many applications where sharing and storing the training samples or their embeddings at each site of deployment is impossible due to privacy, legal, or proprietary constraints. Because of this, most other methods do not compare the test-samples against the training data [13, 21] or make indirect comparisons using auxiliary representations of training data, such as via summary statistics [22], prototypes [29, 30] or synthetic samples [7]. Such derived representations, however, do not faithfully capture all intricacies of the original training data. Hypothesizing that direct comparison with the original training data would be useful to infer whether a test sample is OOD, this paper addresses the following question:
@@
This paper describes a novel OOD detection framework (Fig. 1) with the following key aspects: (1) We introduce Isolation Networks for OOD detection. To infer whether a new test sample is OOD, a neural network is trained to learn a binary classification boundary that separates (isolates) a test sample from the original ID training samples. The network’s convergence rate is then used as a measure of whether the target sample is OOD or similar to the training data. The intuition is that test samples with OOD patterns will be easier to separate from the training samples than ID test samples without OOD patterns. Isolation Networks draw inspiration from Isolation Forests [25], which train decision trees to isolate each training sample. To infer whether a test-sample is OOD, the number of split nodes needed to isolate it is used (c.f. Sec. 2). Isolation Networks instead train a different network for each test sample, to separate it from the training data, and measure convergence as OOD score. (2) We then introduce Decentralized Isolation Networks (DIsoN), a decentralized training framework that enables training Isolation Networks in the practical setting when the ID training data are held on a different computational node than the node of deployment, where the test samples are processed. (3) We extend DIsoN to class-conditional training, where the class of a test sample is first predicted and then used to compare it only against samples in the ID training database of the same class. This reduces variability within the distribution of training samples that the model uses for comparison, making it harder to isolate ID samples that closely resemble their class than OOD samples, improving OOD detection.
@@ 2 Related Work
Federated Learning and OOD Detection. Training of DIsoN uses decentralised optimization similar to Federated Learning (FL). Using FL for OOD detection is a recent area. The few related works, such as  [44, 24], studied settings significantly different from ours: they assume the training data is decentralized across multiple computational nodes, where each node’s data follow a different distribution. Using FL, they train a model that measures distribution-shifts between nodes. After training, the model is applied on test-samples for OOD detection. Our algorithm does not require training data from multiple distributions or multiple nodes. It assumes one node holds all ID training data and a second node holds a single test-sample, which we infer whether it is OOD via decentralised training of a binary classifier. Thus the motivation, use-case and technical challenges are distinct. Moreover, DIsoN may resemble FL with a few-shot client  [33, 38]. Few-shot methods are designed to regularize against overfitting the limited data of few-shot client(s), to train a model for a predictive task (e.g. disease diagnosis) that generalizes to new samples. DIsoN does not train a model for generalization. It optimizes for distinguishing data of the source node from a target sample, using convergence speed to infer if the latter is OOD. Hence, few-shot methods are not directly applicable.
@@ 3.1 Isolation Networks
The core idea of our approach is that the difficulty of training a binary classifier to separate a single target sample $\mathbf{x}_{t}$ from the source data $\mathcal{D}_{s}$ gives an indication whether $\mathbf{x}_{t}$ is ID or OOD. Intuitively, a target sample that is ID will share characteristics with samples in $\mathcal{D}_{s}$. As a result, it is harder for a classifier to isolate (separate) the target sample and will require more update steps. On the other hand, an OOD target sample has patterns that differ from $\mathcal{D}_{s}$, making it easier to isolate and requiring fewer update steps during training. Fig. 2 gives a visual intuition of the idea.
Formally, we consider a neural network consisting of a feature extractor $f:\mathcal{X}\to\mathcal{Z}$, parameterized by $\theta^{f}$, and a binary classification head $h:\mathcal{Z}\to[0,1]$, with parameters $\theta^{h}$. The full network is parameterized by $\theta=(\theta^{f},\theta^{h})$. We use $m\in\{0,1\}$ for isolation labels and reserve $y$ for class labels. In an idealized centralized setting with full access to $\mathcal{D}_{s}$ and the target sample $\mathbf{x}_{t}$, we can train the network $h\circ f$ for the following binary classification problem: Assign label $m=0$ to all source samples $\mathbf{x}_{s}\in\mathcal{D}_{s}$ and label $m=1$ to the target sample $\mathbf{x}_{t}$. Typically, $\mathcal{D}_{s}$ has a high number of samples compared to the single target sample $\mathbf{x}_{t}$, $|\mathcal{D}_{s}|\gg 1$. Therefore, directly training on this highly imbalanced setup is suboptimal, as it can lead to a trivial classifier that always predicts the majority class ($m=0$), ignoring the minority class ($m=1$). To ensure that the target sample has an impact during training, we over-sample $\mathbf{x}_{t}$ within each mini-batch. Specifically, we construct mini-batches as $B=B_{s}\cup\{\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}\}$, where $B_{s}$ is a set of source samples drawn from $\mathcal{D}_{s}$, and $\{\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}\}$ are $N$ replicated copies of the same target sample $\mathbf{x}_{t}$. The empirical loss for this optimization objective is:
$\mathcal{L}_{c}(\theta;B_{s},\mathbf{x}_{t},N)=\frac{1}{|B_{s}|+N}\left(\sum_{\mathbf{x}_{s}\in B_{s}}L(\theta;\mathbf{x}_{s},m=0)+N\cdot L(\theta;\mathbf{x}_{t},m=1)\right),$ (1)
where $L(\theta;x,m)$ denotes the binary cross-entropy loss. In Sec. 3.2 we compare the gradient of the centralized version $g_{c}(\theta;B_{s},\mathbf{x}_{t},N)=\nabla_{\theta}\mathcal{L}_{c}(\theta;B_{s},\mathbf{x}_{t},N)$ to our proposed decentralized version.
@@
Let $p^{(k)}(x)=h(f(x;\theta^{(k)}))$ be the sigmoid‐score after $k$ updates. Fix a window $E_{\mathrm{stab}}\in\mathbb{N}$ and accuracy threshold $\tau\in(0,1]$. The convergence time $K$ is the smallest $k\geq E_{\mathrm{stab}}$ such that
@@
Intuitively, the conditions ensure that the classifier correctly classifies $\mathbf{x}_{t}$ for the last $E_{\mathrm{stab}}$ consecutive updates (first condition), while achieving an accuracy of at least $\tau$ on the source data (second condition). We set $E_{\mathrm{stab}}=5$ and $\tau=0.85$. Further details on these parameter choices are provided in the Appendix D. We define our OOD scoring function as $S(\mathbf{x}_{t})=K$, following the convention where ID samples have higher scores. We emphasize that this setup corresponds to the centralized version of the algorithm, where both $D_{s}$ and $x_{t}$ are accessible on the same node. In Sec. 3.2, we extend it to the decentralized case, where source data and the target sample reside on separate nodes.
@@ 3.2 Decentralized Isolation Networks (DIsoN)
where $g_{S}(\theta)=\frac{1}{|B_{s}|}\sum_{\mathbf{x}_{s}\in B_{s}}\nabla_{\theta}L(\theta;\mathbf{x}_{s},0)$, $g_{T}(\theta)=\nabla_{\theta}L(\theta;\mathbf{x}_{t},1)$ and $N$ is the number of times $\mathbf{x}_{t}$ is oversampled in the centralized version.
@@
Prop. 3.1 shows that our decentralized training exactly replicates the centralized version when $E=1$ and aggregation weights are chosen accordingly. This theoretical equivalence motivated our design of DIsoN: it preserves the core idea of the Isolation Network, while meeting our decentralized data sharing constraints. In practice, we allow $E>1$ to reduce communication overhead, which causes the decentralized updates to deviate from the centralized version. However, as we show in Sec. 4, DIsoN achieves promising results even with the approximation. We found DIsoN to be robust across a broad range of $\alpha$ values, as shown in an ablation study in Sec. 4.2.
Practical Techniques: Augmentation and Normalization. One challenge in DIsoN is to avoid rapid memorization and overfitting of superficial features in the target sample $\mathbf{x}_{t}$, making the isolation task trivial regardless whether the sample is OOD. To prevent this, we apply stochastic data augmentations (e.g. random crops, horizontal flips), which regularize the model to learn invariant features, so that the separation is based on semantic characteristics. Furthermore, in DIsoN, we use Instance Normalization (IN) [35] instead of the widely used Batch Normalization (BN) [16] layers for feature normalization. BN relies on batch-level statistics, which are not suitable for our single-sample TN. IN instead normalizes each feature map per sample and therefore fits better for our decentralized, single target sample scenario.
@@ 4 Experiments & Results
Datasets. We evaluate DIsoN on four publicly available medical imaging benchmark datasets covering dermatology, breast ultrasound, chest X-ray, and histopathology. All datasets consist of real, clinically acquired images and no synthetic data is used. The first three benchmarks use images with naturally occurring non-diagnostic artifacts as OOD samples (e.g., rulers, pacemakers, annotations), while histopathology focuses on semantic and covariate shifts across domains. Example images are shown in Fig.3.
@@
Figure 3: Examples of data. X-Ray: ID X-ray vs. OOD scan with pacemaker. Dermatology: ID lesion vs. OOD image with ruler. Ultrasound: ID artifact-free ultrasound vs. OOD scan containing annotations. Histopathology: ID mitotic‐cell patch vs. near‐OOD patch with different cancer type and far‐OOD patch with different staining.
Training Details. We use a ResNet18 [12] with Instance Normalization (as per Sec.3.2) pre-trained on the dataset-specific task as initialization for DIsoN. DIsoN is trained with Adam (lr=0.001 for dermatology and ultrasound; 0.003 for X-ray). For histopathology SGD with momentum (lr=0.01, momentum=0.9) is used (since [11] suggests pretraining with SGD). Local iterations per communication round are chosen to approximately match one epoch on the training data. We use standard augmentations (e.g. random cropping, rotation, color-jitter) and the aggregation weight is fixed to $\alpha=0.8$, since it performs consistently well across all our experiments (see Sec.  4.2 for effect of $\alpha$). Experiments were run on an NVIDIA RTX A5000. More training details and hyperparameters are provided in the Appendix. OOD detection performance is evaluated with two metrics: (i) area under the receiver operating characteristics curve (AUROC, higher is better) (ii) the false positive rate at 95% true positive rate (FPR95, lower is better). All results in Tables 1 and 2 are averaged over three runs with different seeds, and we report mean and standard deviation. Standard deviations for Tab.2 are in Appendix I due to space limitations.
Table 1: OOD detection performance evaluated on three OOD datasets: Chest X-Ray, Dermatology, and Breast Ultrasound and reported as mean $\pm$ standard deviation over three random seeds. $\downarrow$ means smaller is better and $\uparrow$ means larger is better. Bold numbers highlight the best results, second best results are underlined.
Method Chest X-Ray Dermatology Breast Ultrasound Average
@@
[…]
@@
MSP 54.1/94.8 47.4/93.4 55.7/89.4 59.0/93.5 54.8/94.7 45.6/95.5 56.1/92.4 53.3/93.4 78.4/52.2 82.7/63.6 80.6/57.9
MDS 67.4/80.0 67.2/79.8 65.7/81.9 61.2/87.4 60.6/87.4 55.5/89.5 51.0/91.6 61.2/85.4 87.1/43.6 86.6/39.6 86.8/41.6
fDBD 57.9/83.0 52.5/81.4 57.5/88.7 60.2/89.0 57.9/85.9 51.2/86.2 55.2/92.2 56.1/86.6 74.6/58.3 82.1/63.7 78.3/61.0
ViM 68.0/79.8 66.7/77.8 66.2/77.7 63.5/86.9 61.9/87.3 55.7/91.0 54.9/93.1 62.4/84.8 92.5/29.1 92.6/27.0 92.6/28.0
iForest 37.9/98.3 38.6/96.9 39.7/97.7 41.6/97.5 39.5/98.0 40.4/97.7 46.4/95.7 40.6/97.4 27.7/99.2 32.3/96.8 30.0/98.0
CIDER 71.4/84.6 65.6/89.0 55.5/91.5 64.2/89.8 57.4/92.2 47.7/96.5 64.1/88.2 60.9/90.2 82.5/77.2 95.4/18.3 89.0/47.8
PALM 73.5/78.1 59.2/90.3 66.3/77.7 64.3/93.6 62.1/90.8 41.8/97.6 61.6/93.6 61.2/88.8 97.0/21.8 99.6/1.5 98.3/11.6
$C$-DIsoN 75.4/78.8 79.5/61.7 72.6/79.7 63.0/89.4 64.0/85.3 70.7/79.9 61.8/92.0 69.6/81.0 98.3/4.8 98.4/4.0 98.3/4.4
Baselines. We compare our method against state-of-the-art OOD detection methods from the two main categories: post-hoc and training-time regularization methods. The post-hoc methods include: MSP [13], MDS [22], fDBD [26], ViM [37], Deep iForest [40]. For training-time regularization methods, we evaluate the recent methods CIDER [30] and PALM [29] (both using contrastive learning). We use MDS for OOD scoring on their learned feature representations. Several baselines already use a form of class-conditioning based on the model’s predicted class: MDS, PALM, and CIDER via Mahalanobis distances to class clusters; ViM uses the maximum logit of the predicted class; MSP via softmax probability; and fDBD via distance to the decision boundary (all implicitly or explicitly conditioned on the predicted class). Only iForest does not. CC-DIsoN similarly relies solely on the predicted class from $M_{pre}$ for class-conditional sampling and uses no ground-truth labels, providing no additional information beyond these methods.
4.1 Evaluation on Medical OOD Benchmarks
Dermatology, Chest X-Ray and Breast Ultrasound.